t Bargain Electronics, it costs $33 per unit ($17 variable and $16 fixed) to make an MP3 player at full capacity that normally sells for $40. A foreign wholesaler offers to buy 4,000 units at $30 each. Bargain Electronics will incur special shipping costs of $2 per unit. Assuming that Bargain Electronics has excess operating capacity, indicate the net income (loss) Bargain Electronics would realize by accepting the special order. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Reject
Order
Accept
Order
Net Income
Increase (Decrease)
Revenues $enter a dollar amount $enter a dollar amount $enter a dollar amount
Costs-Manufacturing enter a dollar amount enter a dollar amount enter a dollar amount
         Shipping enter a dollar amount enter a dollar amount enter a dollar amount
Net income $enter a dollar amount $enter a dollar amount $enter a dollar amount
The special order should be select an option                                                          rejected / accepted.
Answer #1

Decision Making

Reject order Accept order Net income increase (decrease)
Revenues 0 4000*30 = 120000 120000
Cost-Manufacturing 0 4000*17 = -68000 -68000
Shipping 0 4000*2 = -8000 -8000
Net income 0 44000 44000

The special order should be accepted